[ I have a 83% in my statistics class and our final is worth 25% of our grade. What do I need to get on the final in order to get at least an 80% in the class? ]
Answer by TimothyLamb(4379)

(100 - 25)(83) + (25)f = 100(80)
(75)(83) + (25)f = 8000
6225 + (25)f = 8000
(25)f = 1775
---
f = 71
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php